Calculate the volume of 0.55M HNO3 necessary to neutralize 5500 cm3 of 0.45M KOH
1
Expert's answer
2015-01-13T11:27:02-0500
For this reaction we can use an equation: M1V1 =M2V2.
So, we get: 0.55*V1 = 0.45*5500; V1 =0.45*5500/0.55 = 4500cm3 of HNO3.